OpenStudy (girlgoyleh.):

Ms. Thomas is making a badge for her son in the shape of a square. The sides of the badge measure 3 inches each. Ms. Thomas wants to put a stripe across the badge from one vertex to the other vertex. What is the approximate length of the stripe?

OpenStudy (girlgoyleh.):

@eliassaab

OpenStudy (anonymous):

\[ \sqrt{ 3^2 + 3^2}=\sqrt{18}=3 \sqrt 2\approx 4.24264 \]